OpenStudy (jaggu):

wht is Integration of 1/(sec x+1)..?

OpenStudy (aravindg):

:)

OpenStudy (aravindg):

ready for integrating?

OpenStudy (aravindg):

shall i start?

OpenStudy (jaggu):

yaa

OpenStudy (jaggu):

r u doing sir

OpenStudy (jaggu):

wht u ..have written ksaimouli..sry I didn't got it...:(

OpenStudy (aravindg):

\[\huge\int\limits \frac{dx}{1+\sec\;x}\] \[\Large =\int\limits \frac{1dx}{1+\frac{1}{\cos x}}\] \[\Large=\int\limits\limits \frac{1dx}{\frac{1+\cos x}{\cos x}}\] \[\Large =\int\limits \frac{\cos xdx}{1+\cos x}\]

OpenStudy (aravindg):

hmm ...do u get any idea frm here?

OpenStudy (jaggu):

yes sir.. after this wht we have to do..?

OpenStudy (aravindg):

try multiplying numerator and denominator by 1-cos x

OpenStudy (jaggu):

okkgot it sir thnxs a lot sir..

OpenStudy (jaggu):

can I ask one more ques..

OpenStudy (aravindg):

:) welcomes!!

OpenStudy (aravindg):

yep bt pls post it separately

OpenStudy (jaggu):

can I ask here only..

OpenStudy (aravindg):

post it separately bcoz long qns laggs the site ,well call me by using @AravindG

OpenStudy (anonymous):

\[\int\limits_{}{1/(1+\sec x)} = x- \tan (x/2)+c\]